# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ 4.1. Entropy method
$+\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+\sigma^{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$ (28)
$+\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\phi_{R}\,{\rm d}x\,{\rm d}t+\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(2\rho_{\varepsilon}+\rho_{\delta}+\phi_{R}\right)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+C\sigma^{\frac{\alpha-1}{\alpha}},$ (29)
where $C>0$ only depends on the $L^{1}$ norm of the initial data. In the last inequality we have used Lemma 1 and the fact that $\int_{\mathbb{R}^{d}}\phi_{R}\,{\rm d}x=C(d)\sigma^{\frac{\alpha-1}{\alpha}}$. From now on, $C$ will be a positive universal constant that can change from line to line.
@@
$-\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\phi_{R}\,{\rm d}x\,{\rm d}t$
@@
$\leq$ $\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t+\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\phi_{R}\,{\rm d}x\,{\rm d}t$
@@
$-\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\sigma C,$ (30)
using Lemma 3 in the last inequality. For the first term in (27c) we get
$\left|\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\phi_{R}\nabla\rho_{\varepsilon}\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|\,{\rm d}x\,{\rm d}t\leq$ $\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}|\log(\rho_{\varepsilon}+\sigma)||\nabla\phi_{R}||\nabla\rho_{\varepsilon}|\,{\rm d}x\,{\rm d}t$
$+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}|\log(\rho_{\delta}+\sigma)||\nabla\phi_{R}||\nabla\rho_{\varepsilon}|\,{\rm d}x\,{\rm d}t.$
Since $|\log(\rho+\sigma)|\leq|\log(\sigma)|+(\rho+\sigma)$ and $|\nabla\phi_{R}|\leq C\sigma^{\frac{1}{\alpha d}}$, we estimate the two integrals above as
$\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}$ $\left|\log(\rho_{\varepsilon}+\sigma)\nabla\phi_{R}\nabla\rho_{\varepsilon}\right|\,{\rm d}x\,{\rm d}t$
$\leq$ $C\sigma^{\frac{1}{\alpha d}+1}\int_{0}^{T}\int_{B_{2R}}\left(|\log(\sigma)||\nabla\rho_{\varepsilon}|+|\nabla\rho_{\varepsilon}|(\rho_{\varepsilon}+\sigma)\right)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;C\sigma^{\frac{1}{\alpha d}+1}|\log(\sigma)|\int_{0}^{T}\int_{B_{2R}}\left(\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}+\rho_{\varepsilon}\right)\,{\rm d}x\,{\rm d}t$
$+\frac{C}{2}\sigma^{\frac{1}{\alpha d}+1}\int_{0}^{T}\int_{B_{2R}}\left(\sigma\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}+\sigma\rho_{\varepsilon}+|\nabla\rho_{\varepsilon}|^{2}+\rho_{\varepsilon}^{2}\right)$
$\leq$ $\;C\sigma,$ (31)
for sufficiently small $\sigma$, by Lemma 1 and Proposition 3 in the last inequality. Similar computations can be done for the integral involving $\rho_{\delta}$.
For the second term in (27c), we compute
$\left|\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\phi_{R}\nabla\rho_{\delta}\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t\right|\leq$ $\;\sigma^{\frac{1}{\alpha d}}\int_{0}^{T}\int_{\mathbb{R}^{d}}|\nabla\rho_{\delta}|(\rho_{\varepsilon}+\sigma)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\;\frac{\sigma^{\frac{1}{\alpha d}}}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(|\nabla\rho_{\delta}|^{2}+\rho_{\varepsilon}^{2}+\sigma\frac{|\nabla\rho_{\delta}|^{2}}{\rho_{\delta}}+\sigma\rho_{\delta}\right)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\sigma^{\frac{1}{\alpha d}}C,$ (32)
@@
using Proposition 1 in the last inequality. Summarizing, by (4.1), (4.1), (4.1) and (4.1), for a sufficiently small $\sigma$, we have the following estimate for the error term
$E_{\varepsilon,\delta}\leq-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+C\,\sigma^{{\frac{1}{\alpha d}}}.$ (33)
Now, we have to estimate the following two terms from (26) involving $\nabla\phi_{R}$:
@@
${I_{1}:=}$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\phi_{R}\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t,$ (34a)
@@
${I_{2}:=}$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\phi_{R}\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$ (34b)
Following the same strategy as in (4.1), for (34a) we get that
$\left|{I_{1}}\right|\lesssim\;C\sigma^{\frac{1}{\alpha d}}|\log(\sigma)|.$
Similarly, (34b) can be estimated as
$\left|{I_{2}}\right|$ $\leq\sigma^{\frac{1}{\alpha d}}\int_{0}^{T}\int_{\mathbb{R}^{d}}|\nabla\rho_{\delta}|\left(1+\frac{|\rho_{\varepsilon}-\rho_{\delta}|}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t$
$\leq\sigma^{\frac{1}{\alpha d}}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\frac{|\nabla\rho_{\delta}|^{2}}{\rho_{\delta}}+\rho_{\delta}+|\nabla\rho_{\delta}|^{2}+\frac{1}{\sigma^{2}}|\rho_{\varepsilon}-\rho_{\delta}|^{2}\right)\,{\rm d}x\,{\rm d}t$
$\leq\sigma^{\frac{1}{\alpha d}}C,$
by a priori estimates and (25).
Now, let us see what happens with the remaining term in (26). These are
${I_{3}}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t,$ (35)
${I_{4}}:=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t.$ (36)
@@
${I_{3}}=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t.$
@@
${I_{3}+I_{4}}=$ $\;-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t$
$+\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t+R_{1}+R_{2},$
@@
$R_{1}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t,$
$R_{2}:=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t.$
@@
$0\leq\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right].$
@@
$R_{1}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\chi\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\delta}|}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\leq\chi\sigma^{1/2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\delta}|}{\sqrt{\rho_{\delta}+\sigma}}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\chi\sigma^{\frac{1}{2}-\frac{1}{2\alpha}}C,$
which vanishes when $\sigma\rightarrow 0$ if $\alpha>1$. We estimate $R_{2}$ in a similar way:
$R_{2}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\frac{\nabla\rho_{\delta}\nabla\rho_{\varepsilon}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\chi\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\leq\chi\sigma^{1/2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|\sqrt{\rho_{\varepsilon}}}{\sqrt{\rho_{\varepsilon}}\sqrt{\rho_{\delta}+\sigma}}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\chi\sigma^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|}{\rho_{\varepsilon}}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\rho_{\varepsilon}-\rho_{\delta}|}{\rho_{\varepsilon}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\delta}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\rho_{\varepsilon}-\rho_{\delta}|^{2}}{\sigma^{2}}\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}}\phi_{R}^{2}\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{1/2}\left(1+\int_{0}^{T}\int_{\mathbb{R}^{d}}(\phi_{R}+\phi_{R}^{2})\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{\frac{1}{2}-\frac{1}{2\alpha}}.$
We conclude that, if $\alpha>1$, $R_{1},R_{2}\rightarrow 0$ when $\sigma\rightarrow 0$.
@@
${I_{3}+I_{4}}=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\times$
×$\times\left(\lvert\nabla\log(\rho_{\varepsilon}+\sigma)\rvert^{2}-\lvert\nabla\log(\rho_{\delta}+\sigma)\rvert^{2}\right)\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$

### v3

@@ 4.1. Contractivity of the relative entropy
$+\delta\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\delta}\frac{\rho_{\varepsilon}-\rho_{\delta}}{\rho_{\delta}+\sigma}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\rho_{\delta})\,{\rm d}x\,{\rm d}t+\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\delta}\,{\rm d}x\,{\rm d}t+\delta\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t$
$\leq$ $2(\varepsilon+\delta)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(\rho_{\varepsilon}+\rho_{\delta}\right)\,{\rm d}x\,{\rm d}t$
$\leq$ ${C(\varepsilon+\delta)},$ (30)
where $C>0$ only depends on the $L^{1}$ norm of the initial data. From now on, $C$ will be a positive universal constant that can change from line to line.
@@
$-\delta\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\,{\rm d}x\,{\rm d}t$
@@
$\leq$ $\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\,{\rm d}x\,{\rm d}t+\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\,{\rm d}x\,{\rm d}t$
@@
$-\delta\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\,{\rm d}x\,{\rm d}t$
$\leq$ $\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t$
$\leq$ ${C(\varepsilon+\delta)},$ (31)
@@
using Lemma 3 in the last inequality. Summarizing, by (4.1) and (4.1) we have the following estimate for the error term
$E_{\varepsilon,\delta}{\leq C(\varepsilon+\delta)}.$ (32)
Now, let us see what happens with the remaining terms in (27). These are
@@
${I_{3}}:=-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t,$ (33)
@@
${I_{4}}:=\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$ (34)
@@
${I_{3}}=$ $-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$
@@
${I_{3}+I_{4}}=$ $\;-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t$
$+\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t+R_{1}+R_{2},$
@@
$R_{1}:=-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t,$
$R_{2}:=\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$
@@
[…]
@@
$R_{1}\leq$ $\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\chi\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{\sigma|\nabla\rho_{\delta}|}{\rho_{\delta}+\sigma}\,{\rm d}x\,{\rm d}t.$
For fixed $\varepsilon,\delta>0$ the expression above tends to zero as $\sigma\to 0$. Indeed, consider the sequence of functions $F_{\sigma}=\frac{\sigma|\nabla\rho_{\delta}|}{\rho_{\delta}+\sigma}$. It holds $F_{\sigma}\leq\overline{F}\equiv|\nabla\rho_{\delta}|=2\sqrt{\rho_{\delta}}|\nabla\sqrt{\rho_{\delta}}|\in L^{2}(0,T;L^{1}(\mathbb{R}^{d}))$. Furthermore $F_{\sigma}\to 0$ a.e. in $\{\rho_{\delta}>0\}$ as $\sigma\to 0$, while $F_{\sigma}=|\nabla\rho_{\delta}|=2\sqrt{\rho_{\delta}}|\nabla\sqrt{\rho_{\delta}}=0$ a.e. in $\{\rho_{\delta}=0\}$. Hence by dominated convergence we obtain that $F_{\sigma}\to 0$ strongly in $L^{1}(0,T;L^{1}(\mathbb{R}^{d}))$. This means
$\limsup_{\sigma\to 0}R_{1}\leq 0\quad\forall\varepsilon,\delta>0.$ (36)
We estimate $R_{2}$ in a similar way:
$R_{2}\leq$ $\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\frac{\nabla\rho_{\delta}\nabla\rho_{\varepsilon}}{\rho_{\delta}+\sigma}\,{\rm d}x\,{\rm d}t{\leq\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}G_{\sigma}\,{\rm d}x\,{\rm d}t,}$
with
$G_{\sigma}\equiv\left[\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{\delta}||\nabla\rho_{\varepsilon}|}{\rho_{\delta}+\sigma}.$
We first observe that (thanks to (35))
$0\leq G_{\sigma}\leq\chi\sigma\frac{|\nabla\rho_{\varepsilon}|}{\rho_{\delta}+\sigma}\leq\chi|\nabla\rho_{\varepsilon}|\equiv\overline{G}$
and $\overline{G}$ is a $\sigma$-independent integrable function. The same estimate also tells us that $G_{\sigma}\to 0$ a.e. on $\{\rho_{\delta}>0\}$ as $\sigma\to 0$. On the other hand, on $\{\rho_{\delta}=0\}$ $G_{\sigma}$ equals a bounded factor (the one between square brackets) times $|\nabla\rho_{\varepsilon}|$ (which is a.e. finite because it is integrable) times $|\nabla\rho_{\delta}|$, which vanishes a.e. in $\{\rho_{\delta}=0\}$ as previously noticed. Again, via the dominated convergence theorem we infer that $G_{\sigma}\to 0$ strongly in $L^{1}(0,T;L^{1}(\mathbb{R}^{d}))$ as $\sigma\to 0$, for fixed (arbitrary) $\varepsilon$, $\delta>0$. Hence we obtain
$\limsup_{\sigma\to 0}R_{2}\leq 0\quad\forall\varepsilon,\delta>0.$ (37)
@@
[…]
$\times\left(\lvert\nabla\log(\rho_{\varepsilon}+\sigma)\rvert^{2}-\lvert\nabla\log(\rho_{\delta}+\sigma)\rvert^{2}\right)\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$